The length of a proposed rectangular flower garden is 4 feet less than five times the width. The area of the proposed garden is 57ft^2. Find the length of the proposed flower garden."

Answer by jorel1380(3719)

w(5w-4)=57
5w2-4w-57=0
(5w-19)(w+3)=0
w=19/5 or -3
Throwing out the negative answer, we get the width to be 19/5 feet and the length to be 15 feet..